pltlaser_day2_ - BS2

REVISION DATE: 10-Feb-2016 20:07:51

This shows the laser data collected by Mike on day 2. Data was collected with only track 8 on. Mike spent some time trying to pin down the 2nd order signal, via laser lines between 339 to 360 nm. The rest of todays lines fill in Kens request for "300-800, every 25 or 50 nm or so". There is still a lot of signal coming from off the UV end of the CCD. I thought the new filter was suppose to take care of that?

Figure 1

image

Figure 2

image

Figure 3

image

Figure 4

image

Figure 5

image

Figure 6

image

Figure 7

image

Figure 8

image

Figure 9

image

Figure 10

image

Figure 11

image

Figure 12

image

Figure 13

image

Figure 14

image

Figure 15

image

Figure 16

image

Figure 17

image

Figure 18

image

Figure 19

image

Figure 20

image

Figure 21

image

Figure 22

image

Figure 23

image

Figure 24

image

Figure 25

image

Figure 26

image

Figure 27

image

Figure 28

image

Figure 29

image

Figure 30

image

Figure 31

image

Figure 32

image

Figure 33

image

Figure 34

image

Figure 35

image

Figure 36

image

Figure 37

image

Figure 38

image

Figure 39

image

pwd: E:\zflora\mldata\mobyrefresh\characterizations\intial\Hawaii-2016-01\BS2cal\day2
Date: 10-Feb-2016 20:07:51
Created from pltlaser_day2_(1)